OpenStudy (anonymous):

The number N of state and federal inmates in millions during year x, where x >2002 can be approximated by the following formula. N=0.03x-58.68 Determine the year in which there were 1.5 million inmates.

OpenStudy (anonymous):

I don't get that all

OpenStudy (anonymous):

set your formula equal to 1.5 and solve for \(x\)

OpenStudy (anonymous):

before we continue, would you like to go back to the previous problem? \[\frac{7}{8}x\geq -\frac{3}{16}\]?

OpenStudy (anonymous):

1.5=0.03x58.68 is that set up right?

OpenStudy (anonymous):

for the previous one, multiply both sides by \(8\) to get \[8\times \frac{7}{8}x\geq -\frac{3}{16}\times 8\] \[\cancel{8}\times \frac{7}{\cancel{8}}x\geq -\frac{3}{\cancel{16}^2}\times \cancel{8}\] \[7x\geq -\frac{3}{2}\]

OpenStudy (anonymous):

then divide both sides by \(7\) to get \[\frac{7x}{7}\geq -\frac{3}{2\times 7}\] \[x\geq -\frac{3}{14}\]

OpenStudy (anonymous):

for this one, solve \[1.5=0.03x-58.68\] for \(x\)

OpenStudy (anonymous):

add \(58.68\) to both sides to get \[1.5+58.68=0.03x\]\[60.18=.03x\]

OpenStudy (anonymous):

then divide both sides by \(0.03\)

OpenStudy (anonymous):

2006 would the year

OpenStudy (anonymous):

after dividing both sides by 0.03

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

now I see how its calculated. thank you!

OpenStudy (anonymous):

yw